# Peer-Aware Cost Estimation in Nonlinear General-Sum Dynamic Games for Mutual Learning and Intent Inference

### v1

@@ III-C Modeling the Learning Dynamics
Remark 2. Notice that in (4), when modeling the policy of the other agent $-k$, the parameter $\theta^{k}$ is included in the policy model. This effectively assumes that the other agent has full knowledge of $\theta^{k}$ and is treated as an expert, which can lead to incorrect and biased estimates as shown in the case studies of this paper.
@@
Remark 3. The proposed learning dynamics $h^{k}_{l}$ follows a simple gradient descent algorithm with a learning rate $\alpha_{k}$, but it is not restricted to gradient-based methods in the N-PACE. It can also take the form of a Bayesian learner or even a neural network model, as explored in [10]. In fact, we implement a Bayesian inference update in all our case studies in Sec. IV, assuming a Gaussian observation model for parameter estimation, to achieve more robust estimation results.
@@
Efficient approximation of the gradient. The learning dynamics proposed in (4) require calculating the gradient of the policy model with respect to the unknown parameter $\hat{\theta}^{-k}_{t}$. The iLQR method, which generates the policy at each step, is an iterative process where the generated trajectories and policy improvement steps depend on the unknown parameter $\hat{\theta}^{-k}_{t}$. While this approach still relates the policy to $\hat{\theta}^{-k}_{t}$, the iterative nature of iLQR can make gradient computation computationally expensive or even intractable for real-time implementation. Notably, in (3), as well as in the converged trajectory $\zeta^{*k}_{t}(\theta^{k},\hat{\theta}^{-k}_{t})$, all parameters are explicitly expressed as functions of $\hat{\theta}^{-k}_{t}$.
@@
To address this challenge, we adopt the approximation method proposed in [34]. Specifically, we ignore the dependency of the converged trajectory in iLQR on $\hat{\theta}^{-k}_{t}$ and replace the gradient of the policy with the gradient of the linear policy obtained by solving the LQ game around the converged trajectory of iLQR. The intuition behind this approximation is that in iLQR, the policy derived from solving the expanded LQ game around the converged trajectory serves as a good approximation of the final converged policy itself, as discussed and demonstrated in [34]. However, automatic differentiation (AD) methods can still be used to compute this gradient if there are no computational constraints. Alternatively, if a different game solver allows the policy to be directly expressed as a function of the unknown parameter $\hat{\theta}^{-k}_{t}$ (e.g., an actor network receiving the parameter as a state input), one can bypass the need for this approximation method.
III-D N-PACE for Multi-Agent Interactions
@@
In Sec. III-C, we proposed the learning dynamics $h^{k}_{l}$ for each agent. However, the update rule in (4) assumes that agent $-k$ is aware of $\theta_{k}$ while generating its policy, effectively treating agent $-k$ as an expert during estimation. This assumption is clearly invalid when all agents are learning. In N-PACE, we address this issue by ensuring that each agent $k$ not only learns its own parameters but also accounts for the learning procedure of the other agent $-k$. This is feasible when agents have access to both learning dynamics, $h^{i}_{l}$ and $h^{j}_{l}$. Consequently, they can replace $\theta_{k}$ with $\hat{\theta}^{k}_{t}$ in equation (4), eliminating the bias in the learning process. As a result, each agent, by having access to learning dynamics and tracking the learning of the other agent, arrives at the following system of learning equations at each step
$\hat{\theta}^{i}_{t+1}$ $=h^{j}_{l}(\hat{\theta}^{i}_{t},\hat{\theta}^{j}_{t},a^{i}_{t},s_{t}),$ (5)
@@
An important property of N-PACE is that when both agents start with the same initial estimates for $\hat{\theta}^{i}_{0}$ and $\hat{\theta}^{j}_{0}$ and continuously track each other’s learning, they will always track the same estimations for both parameters. As a result, the learning dynamics (5) is effectively centralized, even though each agent performs the learning process independently. The assumption of agreeing on initial estimates is not restrictive in many multi-agent interaction scenarios. This holds true when there is a strong prior on the parameters (e.g., a predefined level of rationality or aggressiveness in autonomous driving) or when agents are designed to coordinate despite having different and initially unknown intentions. Even if the initial estimates or learning dynamics of each agent are uncertain (such as human-robot interactions), we refer the readers to works such as [10], which focus on modeling these learning processes and initial estimates in real time, or offline data-driven approaches for modeling the learning dynamics. Alg. 1 summarizes all the steps needed for N-PACE performed by agent $i$.
@@ 3. Learning:
       the learning dynamics as: $\hat{\theta}^{i}_{t+1}=h^{j}_{l}(\hat{\theta}^{i}_{t},\hat{\theta}^{j}_{t},a^{i}_{t},s_{t})$
@@
III-E N-PACE for Intent Communication
The N-PACE algorithm also enables intent communication. Notice that the learning dynamics of agent $-k$, denoted as $h^{-k}_{l}$, are fundamentally conditioned on the actions $a^{k}_{t}$ taken by agent $k$. As a result, each agent can modify its actions to influence the learning process of the other. To enable this, we propose to introduce an additional intent communication term in the form of $\eta(\hat{\theta}^{k}_{t}-\theta^{k})^{2}$ to the instantaneous cost function of each agent, $g^{k}_{\tau}(s_{\tau},a^{i}_{\tau},a^{j}_{\tau};\theta^{k})$, where $\eta$ is the weighting scalar. This leads to the following optimization problem for each agent $k$:
$\min_{a^{k}_{t}}$ $J^{*k}_{t}(s_{t},a^{k}_{t},a^{-k}_{t})=\sum_{\tau=t}^{T}g^{k}_{\tau}(s_{\tau},a^{k}_{\tau},a^{-k}_{\tau};\theta^{k})+\eta\|\hat{\theta}^{k}_{\tau}-\theta^{k}\|^{2}$ (6)
@@ III-E N-PACE for Intent Communication
The influence problem above is computationally expensive and may be intractable for real-time implementation. Remember that the learning dynamics $h^{k}_{l},h^{-k}_{l}$ themselves require solving a general-sum game, using iLQR in this case, for their prediction phase and computing the gradient of that prediction. To address this challenge, we propose an approximation method for solving the optimization problem in (6). Instead of considering the evolution of the other agent’s parameter learning throughout the entire horizon $[t,T]$ in (6) to form the cumulative cost $J^{*k}$, we focus only on a single-step update of the other agent’s learning as $\hat{\theta}^{-k}_{{t}+1}=h^{k}_{l}(\hat{\theta}^{-k}_{t},\hat{\theta}^{k}_{t},a^{-k}_{t},s_{t})$, so we can modify the cumulative cost of agent $k$ at each step as
@@
where $\hat{\theta}^{k}_{t+1}$ depends on $a^{k}_{t}$ through the learning dynamics $h^{-k}_{l}$. We will show the effectiveness of this approximation for intent communication in Case Study 3 in Sec. IV.
IV Case Studies
In this section, we demonstrate the importance of modeling the learning dynamics of the other agent using the proposed N-PACE for task completion, safety, and signaling in multi-agent systems. We illustrate this through three simulation case studies: 1) an assistive nonlinear lunar lander, 2) an autonomous car in lane merging, and 3) an autonomous car in intersection crossing. The learning dynamics in all case studies employ gradient descent-based Bayesian inference using a Gaussian belief.
IV-A Case Study 1: Assistive Lunar Lander
In this example, we study a shared control problem to show the importance of N-PACE in task completion. An assistive lunar lander scenario in which a human agent controls only the angle of the lunar lander $\phi$ by applying the torque $T$, while an autonomous agent controls the thrust force $F$ along the direction of the lunar lander’s angle (assuming no gravity). Both agents cooperate to land the lunar lander at the final position $(x_{f},y_{f})$, but with asymmetric information: the human knows only the final desired $x$-position $x_{f}$, while the autonomous agent knows only the final desired $y$-position $y_{f}$. As a result, the human must learn the parameter $\hat{y}_{f}$, and the autonomous agent must learn $\hat{x}_{f}$ based on observed interactions. The underlying nonlinear dynamics and the running costs for the human $g_{h}$ and the agent $g_{a}$ are defined as:
$\dot{x}$ $=v_{x},\quad\dot{y}=v_{y},\quad\dot{\phi}=\omega,\quad\dot{v}_{x}=F\sin{\phi},\quad\dot{v}_{y}=F\cos{\phi},$
$\dot{\omega}$ $=T,\quad g_{a}=10(x-\hat{x}_{f})^{2}+10(y-y_{f})^{2}+10(\phi-\frac{\pi}{2})^{2}$
$+10(v_{x}^{2}+v_{y}^{2}+\omega^{2})+F^{2},\quad g_{h}=10(x-x_{f})^{2}+$
$10(y-\hat{y}_{f})^{2}+10(\phi-\frac{\pi}{2})^{2}+10(v_{x}^{2}+v_{y}^{2}+\omega^{2})+T^{2}.$
The game has a time horizon of $T=5s$ with a sampling time of $\Delta T=0.1s$ that is used to discretize the above dynamics. In this example, we set $x_{f}=5$ and $y_{f}=-5$, while the lunar lander starts at the initial position $(-5,8)$ with a zero initial angle. We conducted three simulations for this case study: 1) a complete information game where both agents have full knowledge of $x_{f}$ and $y_{f}$ as ground truth, 2) a scenario where each agent treats its peer as an expert, assuming the other agent has complete information and 3) a scenario in which agents operate according to N-PACE.
Figure 2 shows summarizes the results, where Fig. 2 (a) presents the trajectory of the lunar lander under different algorithms, explicitly demonstrating the superiority of N-PACE compared to the approach where each agent treats its peer as an expert. It can be seen that the final position of the lunar lander in N-PACE is closer to the desired final position, and the entire trajectory aligns more closely with the complete information game trajectory (the local Nash equilibrium). In contrast, treating the other agent as an expert results in the lunar lander landing at a different position, underscoring the importance of modeling the learning dynamics of the other agent for successful task completion. Figs. 2 (b) and (c) present the results of parameter learning performance for each agent under different algorithms, illustrating how N-PACE successfully converges to the true final position values for each agent, whereas treating the other agent as an expert results in biased parameter learning. Notably, treating the other agent as an expert can lead to a significant overshoot in the learning process. Intuitively, this may be due to a high initial bias in assuming that the peer agent has complete information.

### v2

@@ 3.2. Modeling the Learning Dynamics
(Expert Peer Assumptions) Notice that in (4), when modeling the policy of the other agent k-k, the parameter $\theta^{k}$ is included in the policy model. This effectively assumes that the other agent has full knowledge of $\theta^{k}$ and is treated as an expert, which can lead to incorrect and biased estimates.
Remark 0.
@@
(Flexibility of Learning Dynamics) The proposed learning dynamics $h^{k}_{l}$ follows a simple gradient descent algorithm with a learning rate $\alpha_{k}$, but it is not restricted to gradient-based methods in the N-PACE. It can also take the form of a Bayesian learner or even a neural network model, as explored in tian2023towards.
@@
Efficient approximation of the gradient. The learning dynamics proposed in (4) require calculating the gradient of the policy model with respect to the unknown parameter $\hat{\theta}^{-k}_{t}$. The ILQ method, which generates the policy at each step, is an iterative process where the generated trajectories and policy improvement steps depend on the unknown parameter $\hat{\theta}^{-k}_{t}$. While this approach still relates the policy to $\hat{\theta}^{-k}_{t}$, the iterative nature of ILQ can make gradient calculation computationally expensive or even intractable for real-time implementation. Notably, in (3), as well as in the converged trajectory $\zeta^{*k}_{t}(\theta^{k},\hat{\theta}^{-k}_{t})$, all parameters are explicitly expressed as functions of $\hat{\theta}^{-k}_{t}$.
@@
To address this challenge, we adopt the approximation method proposed in li2023cost. Specifically, we ignore the dependency of the converged trajectory in ILQ on $\hat{\theta}^{-k}_{t}$ and replace the gradient of the policy with the gradient of the linear policy obtained by solving the LQ game around the converged trajectory of iLQR. The intuition behind this approximation is that in ILQ, the policy derived from solving the expanded LQ game around the converged trajectory serves as a good approximation of the final converged policy itself, as discussed and demonstrated in li2023cost. However, automatic differentiation (AD) methods can still be used to compute this gradient if there are no computational constraints. Alternatively, if a different game solver allows the policy to be directly expressed as a function of the unknown parameter $\hat{\theta}^{-k}_{t}$ (e.g., an actor network receiving the parameter as a state input), one can bypass the need for this approximation method.
3.3. N-PACE for Multi-Agent Interactions
@@
In Sec. 3.2, we proposed the learning dynamics $h^{k}_{l}$ for each agent. However, the update rule in (4) assumes that agent k-k is aware of $\theta_{k}$ while generating its policy, effectively treating agent k-k as an expert during estimation. This assumption is clearly invalid when all agents are learning. In N-PACE, we address this issue by ensuring that each agent $k$ not only learns parameters of the other agent, but also accounts for the learning procedure of the other agent k-k. This is feasible when agents have access to both learning dynamics, $h^{i}_{l}$ and $h^{j}_{l}$. Consequently, they can replace $\theta_{k}$ with $\hat{\theta}^{k}_{t}$ in equation (4), eliminating the bias in the learning process. As a result, each agent, by having access to learning dynamics and tracking the learning of the other agent, arrives at the following system of learning equations at each step
$\hat{\theta}^{i}_{t+1}$ $=h^{i}_{l}(\hat{\theta}^{i}_{t},\hat{\theta}^{j}_{t},a^{i}_{t},s_{t}),$ (5)
@@
Details of N-PACE are shown in Algorithm 1. It should be noted that N-PACE does not require both agents to follow the same learning dynamics; what matters in N-PACE is that agents have knowledge of each other’s learning dynamics. An important property of N-PACE is that when both agents start with the same initial estimates for $\hat{\theta}^{i}_{0}$ and $\hat{\theta}^{j}_{0}$ and continuously track each other’s learning, they will always track the same estimations for both parameters. As a result, the learning dynamics (5) are effectively centralized, even though each agent performs the learning process independently. The situation where agents know others’ initial priors/estimates often happens in multi-agent settings when there is a strong prior on the parameters (e.g., a predefined level of rationality or aggressiveness in autonomous driving) or when agents are designed to coordinate despite having different unknown intentions.
Remark 0.
(Unknown learning dynamics) If the initial estimates or learning dynamics of each agent are uncertain (such as human-robot interactions), we refer the readers to works such as tian2023towards; xie2021learning; wang2022influencing, which focus on data-driven modeling of these learning processes or latent representation of such learning behaviors.
Even if agents misspecify each other’s initial estimates, N-PACE tends to attenuate the resulting bias compared to modeling the peer as a complete information agent that introduces a fixed bias into the learning procedure. Intuitively, for example, when the peer’s update map admits a stable fixed point, the coupled updates steer the estimates toward that fixed point, reducing the initial mismatch over time, trying to keep N-PACE robust to the initial mismatch. In contrast, methods that approximate the peer as a fully informed “expert” introduce a structural bias: the model mismatch is baked into the inference mechanism and therefore persists throughout the interaction liu2016blame; soltanian2025pace. A robustness analysis of N-PACE to misspecified learning dynamics and initial estimates has been provided in Sec 4.2. While N-PACE’s advantage in eliminating the inference bias of baseline methods is intuitive, the problem it addresses is generally non-convex; thus, any convergence guarantee necessarily depends on the stage-cost structure and the underlying dynamics. But to make the advantage theoretically more precise—relative to baselines that treat the peer as an “expert”—we analyze a class of dynamic games in which each agent’s policy is linear with respect to its own intent parameter $\theta^{k}$. This structure arises when the Hessian of the stage cost $g_{k}(\cdot)$ with respect to states and control actions is independent of $\theta^{k}$, while the gradient terms depend linearly on $\theta^{k}$. Such a setting is common in goal-reaching tasks where the intent parameter directly represents the goal position.
Proposition 0 (N-PACE v.s treating the peer as an expert).
Suppose the stage costs $g_{t}^{k}(s_{t},a_{t}^{i},a_{t}^{j};\theta^{k})$ in (2) have Hessians with respect to $(s,a^{i},a^{j})$ that are independent of $\theta^{k}$, with $\theta^{k}$ entering only linearly in the first-order terms, and on any compact operating set $s_{t}\in\Omega$, the policy of each agent admits the form $\pi_{t}^{k}(s_{t};\theta^{k},\hat{\theta}_{t}^{-k})=M_{t}^{k}(s_{t},\hat{\theta}_{t}^{-k})\,\theta^{k}+S_{t}^{k}(s_{t},\hat{\theta}_{t}^{-k})$ with $M_{t}^{k},S_{t}^{k}$ bounded. Also, suppose each agent updates its peer parameter online by the gradient descent learning rule (4), then the convergence of intent estimation in N-PACE is guaranteed for sufficiently small learning rate $\alpha$, whereas methods that treat the peer as an expert/complete-information agent can fail to converge for any value of learning rate.
Proof.
See Appendix A. ∎
@@ 3. Learning:
    the learning dynamics as: $\hat{\theta}^{i}_{t+1}=h^{i}_{l}(\hat{\theta}^{i}_{t},\hat{\theta}^{j}_{t},a^{i}_{t},s_{t})$
@@
3.4. N-PACE for Intent Communication
The N-PACE algorithm also enables intent communication. Notice that the learning dynamics of agent k-k, denoted as $h^{-k}_{l}$, are fundamentally conditioned on the actions $a^{k}_{t}$ taken by agent $k$. As a result, each agent can modify its actions to influence the learning process of the other. To enable this, we propose to introduce an additional intent communication term in the form of $\eta(\hat{\theta}^{k}_{t}-\theta^{k})^{2}$ to the instantaneous cost function of each agent, $g^{k}_{\tau}(s_{\tau},a^{i}_{\tau},a^{j}_{\tau};\theta^{k})$, where $\eta$ is the weighting scalar. This leads to the following optimization problem for each agent $k$:
$\min_{a^{k}_{t}}$ $J^{*k}_{t}(s_{t},a^{k}_{t},a^{-k}_{t})=\sum_{\tau=t}^{T}\Big[\,g^{k}_{\tau}(s_{\tau},a^{k}_{\tau},a^{-k}_{\tau};\theta^{k})+\eta\,\|\hat{\theta}^{k}_{\tau}-\theta^{k}\|^{2}\,\Big]$ (6)
@@ 3.4. N-PACE for Intent Communication
The influence problem above is computationally expensive and may be intractable for real-time implementation. Remember that the learning dynamics $h^{k}_{l},h^{-k}_{l}$ themselves require solving a general-sum game, using ILQgames in this case, for their prediction phase and computing the gradient of that prediction. To address this challenge, we propose an approximation method for solving the optimization problem in (6). Instead of considering the evolution of the other agent’s parameter learning throughout the entire horizon $[t,T]$ in (6) to form the cumulative cost $J^{*k}$, we focus only on a single-step update of the other agent’s learning as $\hat{\theta}^{-k}_{{t}+1}=h^{k}_{l}(\hat{\theta}^{-k}_{t},\hat{\theta}^{k}_{t},a^{-k}_{t},s_{t})$, so we can modify the cumulative cost of agent $k$ at each step as
@@
where $\hat{\theta}^{k}_{t+1}$ depends on $a^{k}_{t}$ through the learning dynamics $h^{-k}_{l}$.